OpenStudy (anonymous):

How do you plug in the following equations into a graphing calculator to get the answer? 3x + 6y = 18 2x + 5y = 20

OpenStudy (amistre64):

y1 = Eq1 y2 = Eq2 hit graph and see where they intersect

OpenStudy (amistre64):

cant really say much more since every calc is different

OpenStudy (anonymous):

yeah but where do I put the =18? u cant do that in my graphing calculator

OpenStudy (amistre64):

ooo ... put it to the other side :)

OpenStudy (anonymous):

how? xD

OpenStudy (amistre64):

y1 = 3x + 6y = 18 y1 = 3x + 6y - 18 y2 = 2x + 5y = 20 y2 = 2x + 5y - 20

OpenStudy (amistre64):

its just on the wrong side of the equals sign; so bring it over like you normally would :)

OpenStudy (anonymous):

oooooh thank u :)

OpenStudy (anonymous):

wait wait u can't put y in the equation since it already has one

OpenStudy (amistre64):

..... fine, solve each equation for y then :)

OpenStudy (amistre64):

3x + 6y1 = 18 y1 = -3x/6 +18/6 = -x/2 +3 2x + 5y2 = 20 y2 = -2x/5 + 20/5 = -2x/5 + 4

OpenStudy (anonymous):

oh thnx aloooooooooooot ^_^